# Familiarity-Aware Evidence Compression for Retrieval-Augmented Generation

### v2

@@ Abstract
Retrieval-augmented generation (RAG) improves large language models (LMs) by incorporating non-parametric knowledge through evidence retrieved from external sources. However, it often struggles to cope with inconsistent and irrelevant information that can distract the LM from its tasks, especially when multiple evidence pieces are required. While compressing the retrieved evidence with a compression model aims to address this issue, the compressed evidence may still be unfamiliar to the target model used for downstream tasks, potentially failing to utilize the evidence effectively. We propose FaviComp (Familiarity-aware Evidence Compression), a novel training-free evidence compression technique that makes retrieved evidence more familiar to the target model, while seamlessly integrating parametric knowledge from the model. Experimental results show that FaviComp consistently outperforms most recent evidence compression baselines across multiple open-domain QA datasets, improving accuracy by up to 28.1% while achieving high compression rates. Additionally, we demonstrate the effective integration of both parametric and non-parametric knowledge during evidence compression [^1] Code and data are available at https://github.com/luka-group/FaviComp.
@@
Dongwon Jung1  Qin Liu1  Tenghao Huang2  Ben Zhou3  Muhao Chen1 1University of California, Davis, 2University of Southern California, 3Arizona State University {dwojung,muhchen}@ucdavis.edutenghaoh@usc.edubenzhou@asu.edu
@@ 1 Introduction
Nevertheless, RAG often struggles to cope with inconsistent and irrelevant information from the multiple evidence set, which can interfere with downstream tasks (Shi et al., 2023). This highlights the need for evidence compression to identify and retain only the essential information for LMs to utilize effectively. Traditionally, evidence compression has focused on reranking documents or sentences by relevance and then incorporating a top-ranked subset (Nogueira et al., 2020; Zhuang et al., 2023; Wang et al., 2023c) or compressing the documents into a compact form that retains only essential context (Jiang et al., 2023a; Xu et al., 2024; Yoon et al., 2024). However, the compressed evidence might be unfamiliar to the LM employed for the downstream task (referred to as the target model), particularly due to discrepancies in the internal knowledge and prompt preferences between the compression model and the target model (Gonen et al., 2023; Lee et al., 2024; Li et al., 2024a; Mallen et al., 2023). When LMs encounter unfamiliar contextual information, they often fail in balancing parametric and non-parametric knowledge, either by overly relying on their parametric knowledge (Longpre et al., 2021; Wang et al., 2023a; Zhou et al., 2023) or by utilizing retrieved evidence without considering its relevance to the input (Wu et al., 2024).
To address these challenges, we propose Familiarity-aware Evidence Compression (FaviComp), a training-free evidence compression method that consolidates multiple evidence into an abstractive summary that is more familiar to the target model, while seamlessly integrating parametric knowledge from the model. Inspired by the prior findings that an LM’s familiarity with a prompt is generally reflected by low perplexity (Liu et al., 2024; Gonen et al., 2023; Wang et al., 2023b), FaviComp proactively composes the compressed evidence in a way to lower the perplexity of the target model. Specifically, instead of directly selecting the highest probability token from the compression model at each decoding step, FaviComp selects the token from the ensemble of the token probabilities from both the compression and target models. This ensemble decoding therefore constrains the token search space of the compression model to those with lower perplexity for the target model, making the context more familiar to the target model (Liu et al., 2024).
@@ 2 Method
We present FaviComp, a decoding-time evidence compression method that familiarizes retrieved evidence with the target model while synergizing them with the model’s parametric knowledge. We first illustrate the motivation for FaviComp in Section 2.1 and provide the preliminaries of evidence compression in RAG Section 2.2, followed by a detailed definition of our proposed framework in Section 2.3.
@@ 2.1 Motivation and Method Overview
Figure 1 illustrates the overview of FaviComp. Existing evidence compression methods employ the compression model to filter out irrelevant information from the retrieved documents. However, since the compression model and the target model are different, the target model might not be familiar to the compressed evidence due to the difference in internal knowledge and prompt preferences between the two models (Gonen et al., 2023; Lee et al., 2024; Mallen et al., 2023). In addition, the compressed evidence cannot be supplemented with the rich parametric knowledge from the target model. In the example, even though the compression model successfully summarizes the essential information, the target model produces an inaccurate answer due to the unfamiliarity with the target model and the lack of integration of the parametric knowledge. On the other hand, FaviComp compresses the given evidence more favorable to the target model by using a novel ensemble decoding technique and leverages its parametric knowledge to supplement the missing evidence (“Lionel Messi made his league debut in Barcelona"), effectively combining evidential and parametric knowledge.
@@ 2.2 RAG with Evidence Compression
Given a set of $k$ retrieved evidence snippets $D=\{d_{1},d_{2},\ldots,d_{k}\}$ and a textual input sequence $x$, RAG aims to generate an output sequence $y$, conditioned on both $D$ and $x$. However, RAG directly utilizes $D$ which often contains irrelevant information to $x$, potentially confusing the target model in downstream tasks (Shi et al., 2023). Thus, we uses an additional compression model to condense $D$ into a concise and input-relevant context $c$, which is then used in place of $D$ during the downstream generation process. Thus, the RAG with evidence compression is formalized as:
$y^{*}=\arg\max_{y}P_{t}(y\mid x,\hat{c}),$
$\hat{c}=P_{c}(c\mid x,[d_{1},d_{2},\ldots,d_{k}]),$
where $y^{*}$ is the final output sequence, [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes concatenation, and $P_{t}$ and $P_{c}$ represent the probability distributions of the target and compression models, respectively. In this work, we consider any natural language prompting tasks, such as open-domain QA tasks, where $x$ represents the input prompt (also known as the query in QA tasks) and $y^{*}$ denotes the output sequence.
@@
$P_{c}(c\mid I_{comp},x,D)=\prod_{i=1}^{|c|}P_{c}(c_{i}|I_{comp},x,D,c_{<i}),$
where $|c|$ is the length of the summary $c$.
@@ 2.3 Ensemble Decoding for FaviComp
Specifically, the target model is directed to generate a context $c$ that would be helpful in answering the question $x$ without referencing the evidence set. This is also done in zero-shot using a context generation instruction prompt $I_{gen}$ such as:
@@
$P_{t}(c\mid I_{gen},x)=\prod_{i=1}^{|c|}P_{t}(c_{i}|I_{gen},x,c_{<i}),$
where $|c|$ denotes the length of the generated context $c$.
@@
$c_{i}$ $=\arg\max_{c_{i}\in V}(\alpha\cdot\log P_{t}(c_{i}|I_{gen},x,c_{<i})$
$+(1-\alpha)\cdot\log P_{c}(c_{i}|I_{comp},x,D,c_{<i})),$
where $c_{i}$ is the subsequent token, and $\alpha$ is the ensemble coefficient that weighs between the two probability distributions. We demonstrate how the coefficient $\alpha$ impacts both the perplexity and the downstream performance in Section 4.2.
Ensemble decoding proactively shifts the token search space in evidence compression by upweighting those tokens with lower perplexity from the target model’s perspective (Liu et al., 2024), resulting in a compressed evidence that is more familiar to the target model. Note that since both objectives ultimately share the goal of generating context relevant to the question, combining the logits ensures alignment with this ultimate goal.
In addition, ensemble decoding enables FaviComp to seamlessly integrate both retrieval knowledge from the external evidence set and the target model’s parametric knowledge. Specifically, FaviComp selects the $\arg\max$ token from the target model only when the token’s probability is higher than that of the compression model, demonstrating that FaviComp draws on parametric knowledge only when necessary—potentially when the compression model is uncertain about the next token. This is particularly beneficial for complex tasks like multi-document QA, where the evidence set may not include all the necessary information (Mallen et al., 2023). In such cases, the missing information in compressed evidence can be supplemented by tokens generated from context generation by the target model, which is entirely based on parametric knowledge. We demonstrate in Section 4.3 and Section 5 that FaviComp can incorporate knowledge from both sources effectively, leading to a performance boost compared to compression methods that solely focus on distilling knowledge from the evidence set.
@@ 3.2 Implementation Details
For all the comparison methods, we utilize Llama3-8B-Instruct and Mixtral-8x7B- Instruct as the target model to tackle downstream QA tasks with RAG. For FaviComp and Zero-shot Summarization, we employ two compression models, one for each target model: Llama3.2-3B-Instruct for Llama3- 8B-Instruct target model and Mistral-7B- Instruct for Mixtral-8x7B-Instruct target model. For each question, we retrieve five documents from 2018 Wikipedia corpus (Karpukhin et al., 2020) using Contriever-MSMARCO (Izacard et al., 2021), so as to be consistent with previous studies (Xu et al., 2024; Yoon et al., 2024). We set ensemble coefficient $\alpha$ of FaviComp to 0.5 by default, for which more analyses are given in Section 4.2. The prompts used in the experiment are presented in Appendix C.
@@ 3.3 Baselines
We consider the following categories of baselines. (1) No Context: RAG without any context. (2) Gold Compression: RAG using directly relevant evidence from the retrieved documents if they exist. (3) Raw Document: RAG with raw documents that have not undergone any compression. (4) Generated Context (Yu et al., 2023): RAG with context generated by the same LM as the target model. This is equivalent to FaviComp with $\alpha=1$, as we rely solely on the target model to generate context when $\alpha=1$. (5) Reranking-based Methods: We rerank sentences in the evidence set and choose top-ranked sentences as the context. We utilize two rerankers—Sentence-BERT (Reimers and Gurevych, 2020) and RECOMP-extractive (Xu et al., 2024). (6) Compression-based Methods: We employ four compressors—LongLLMLingua (Jiang et al., 2023a), RECOMP-abstractive (Xu et al., 2024), CompAct (Yoon et al., 2024), and Zero-shot Summarization. For Zero-shot Summarization, we use the same evidence compression instruction prompt of FaviComp to summarize multiple evidence using the same LM as the target model. This is equivalent to FaviComp with $\alpha=0$, as we depend entirely on the compression model without any intervention from the target model. A detailed explanation of the implementation of the baselines is provided in Appendix A.
@@
Table 1: Experimental results on five open-domain QA datasets. Size column represents the size of the compression model used for each method. indicates a fully-supervised compression model, where the compressor is trained. The best Accuracy and token-level F1 scores for each dataset are in bold.
@@ 4 Experimental Results
In this section, we compare the overall performance of FaviComp with other baselines across the five datasets (Section 4.1), explore the impact of ensemble coefficient $\alpha$ on performance and perplexity (Section 4.2), investigate how effectively FaviComp incorporate parametric and non-parametric knowledge (Section 4.3), and compare the compression rates with other baselines (Section 4.4).
@@ 4.1 Main Results
The overall performance of FaviComp and the baselines across the five datasets are presented in Table 1.222We present additional experimental results using other combinations of compression and target model at Section B.1. To start with, the compression-based methods consistently outperform the reranking-based methods, due to the fact that the reranking-based methods are prone to losing more question-relevant information by discarding lower-ranked sentences.
@@
Finally, given that Zero-shot Summarization corresponds to FaviComp with $\alpha=0$ and Generated Context corresponds to FaviComp with $\alpha=1$, the fact that FaviComp outperforms both baselines highlights its ability to effectively incorporate tokens from both sources—evidence summary and generated context. This results in superior performance compared to relying on just one source alone.
@@ 4.2 Impact of Ensemble Coefficient on Performance and Perplexity
Figure 2 illustrates how performance and perplexity change as the ensemble coefficient $\alpha$ is varied across the values when using Llama3.2-3B-Instruct and Llama3-8B- Instruct compression-target pairs on HQA and MQ datasets [^4] Results for other datasets are included in Figure 6.. We calculate the perplexity of the compressed evidence conditioned on the preceding inputs, i.e. instruction, demonstrations, and the question. For all the datasets, performance is the highest when $\alpha=0.5$, indicating that proactively lowering perplexity by equally weighting both input sources yields the best results. When $\alpha$ is below 0.5, performance improves as the perplexity of compressed evidence decreases, which aligns with the previous works (Liu et al., 2024; Gonen et al., 2023). However, when $\alpha$ exceeds 0.5, performance declines as perplexity decreases due to the lack of evidential knowledge during evidence compression. Additionally, when $\alpha$ reaches 0.9 or 1.0, there is a slight rise in the perplexity due to LM’s increased uncertainty with limited evidential knowledge.

### v3

@@ Abstract
Retrieval-augmented generation (RAG) improves large language models (LMs) by incorporating non-parametric knowledge through evidence retrieved from external sources. However, it often struggles to cope with inconsistent and irrelevant information that can distract the LM from its tasks, especially when multiple evidence pieces are required. While compressing the retrieved evidence with a compression model aims to address this issue, the compressed evidence may still be unfamiliar to the target model used for downstream tasks, potentially failing to utilize the evidence effectively. We propose FaviComp (Familiarity-aware Evidence Compression), a novel inference-time evidence compression technique that makes retrieved evidence more familiar to the target model, while seamlessly integrating parametric knowledge from the model. Experimental results show that FaviComp consistently outperforms the most recent evidence compression baselines across multiple open-domain QA datasets, improving accuracy by up to 28.1% while achieving high compression rates. Additionally, we demonstrate the effective integration of both parametric and non-parametric knowledge during evidence compression.  [^1] Code and data are available at https://github.com/luka-group/FaviComp
@@
Dongwon Jung1   Qin Liu1   Tenghao Huang2   Ben Zhou3   Muhao Chen1 1University of California, Davis, 2University of Southern California, 3Arizona State University {dwojung,qinli,muhchen}@ucdavis.edutenghaoh@usc.edubenzhou@asu.edu
@@ 1 Introduction
Nevertheless, RAG often struggles to cope with inconsistent and irrelevant information from the multiple evidence pieces, which can interfere with downstream tasks (Shi et al., 2023). This highlights the need for evidence compression to identify and retain only the essential information for LMs to utilize effectively. Traditionally, evidence compression has focused on reranking documents or sentences by relevance and then incorporating a top-ranked subset (Nogueira et al., 2020; Zhuang et al., 2023; Wang et al., 2023c) or compressing the documents into a compact form that retains only essential context (Jiang et al., 2023a; Xu et al., 2024; Yoon et al., 2024). However, the compressed evidence might be unfamiliar to the LM employed for the downstream task (referred to as the target model), particularly due to discrepancies in the internal knowledge and prompt preferences between the compression model and the target model (Gonen et al., 2023; Lee et al., 2024; Li et al., 2024a; Mallen et al., 2023). When LMs encounter unfamiliar contextual information, they often fail in balancing parametric and non-parametric knowledge, either by overly relying on their parametric knowledge (Longpre et al., 2021; Wang et al., 2023a; Zhou et al., 2023) or by utilizing retrieved evidence without considering its relevance to the input (Wu et al., 2024).
To address these challenges, we propose Familiarity-aware Evidence Compression (FaviComp), an inference-time evidence compression method that consolidates multiple evidence into an abstractive summary that is more familiar to the target model, while seamlessly integrating parametric knowledge from the model. Inspired by the prior findings that an LM’s familiarity with a prompt is generally reflected by low perplexity (Liu et al., 2024; Gonen et al., 2023; Wang et al., 2023b), FaviComp proactively composes the compressed evidence in a way to lower the perplexity of the target model. Specifically, instead of directly selecting the highest probability token from the compression model at each decoding step, FaviComp selects the token from the ensemble of the token probabilities from both the compression and target models. This ensemble decoding therefore constrains the token search space of the compression model to those with lower perplexity for the target model, making the context more familiar to the target model (Liu et al., 2024).
@@ 2 Method
We present FaviComp, a inference-time evidence compression method that familiarizes retrieved evidence with the target model while synergizing them with the model’s parametric knowledge. We first illustrate the motivation for FaviComp in Section˜2.1 and provide the preliminaries of evidence compression in RAG Section˜2.2, followed by a detailed definition of our proposed framework in Section˜2.3.
@@ 2.1 Motivation and Method Overview
Figure˜1 illustrates the overview of FaviComp. Existing evidence compression methods employ the compression model to filter out irrelevant information from the retrieved documents. However, since the compression model and the target model are different, the target model might not be familiar to the compressed evidence due to the difference in internal knowledge and prompt preferences between the two models (Gonen et al., 2023; Lee et al., 2024; Mallen et al., 2023). In addition, the compressed evidence cannot be supplemented with the rich parametric knowledge from the target model. In the example, even though the compression model successfully summarizes the essential information, the target model produces an inaccurate answer due to the unfamiliarity with the target model and the lack of integration of the parametric knowledge. On the other hand, FaviComp compresses the given evidence more favorable to the target model by using a novel ensemble decoding technique and leverages its parametric knowledge to supplement the missing evidence (“Lionel Messi made his league debut in Barcelona"), effectively combining evidential and parametric knowledge.
@@ 2.2 RAG with Evidence Compression
Given a set of $k$ retrieved evidence snippets $D=\{d_{1},d_{2},\ldots,d_{k}\}$ and a textual input sequence $x$, RAG aims to generate an output sequence $y$, conditioned on both $D$ and $x$. However, RAG directly utilizes $D$ which often contains irrelevant information to $x$, potentially confusing the target model in downstream tasks (Shi et al., 2023). Thus, we use an additional compression model to condense $D$ into a concise and input-relevant context $c$, which is then used in place of $D$ during the downstream generation process. Thus, the RAG with evidence compression is formalized as:
[…]
$\hat{c}=P_{\text{comp}}(c\mid x,[d_{1},d_{2},\ldots,d_{k}]),$
where $y^{*}$ is the final output sequence, $\cdot$$\cdot$] denotes concatenation, and $P_{\text{tar}}$ and $P_{\text{comp}}$ represent the probability distributions of the target and compression models, respectively. In this work, we consider any natural language prompting tasks, such as open-domain QA tasks, where $x$ represents the input prompt (also known as the query in QA tasks) and $y^{*}$ denotes the output sequence.
@@
$P_{\text{comp}}(c\mid\mathcal{C}_{\text{comp}})=\prod_{i=1}^{|c|}P_{\text{comp}}(c_{i}\mid\mathcal{C}_{\text{comp}},c_{<i}),$
where $\mathcal{C}$ denotes the input prompt, constructed by stringifying $\{I_{\text{comp}},x,D\}$ using a predefined prompt template and $|c|$ is the length of the summary $c$.
@@ 2.3 Ensemble Decoding for FaviComp
Specifically, the target model is instructed to generate a context $c$ that would be helpful in answering the question $x$ without referencing the evidence set. This is also done in zero-shot using a context generation instruction prompt $I_{gen}$ such as:
@@
$P_{\text{tar}}(c\mid\mathcal{C}_{\text{gen}})=\prod_{i=1}^{|c|}P_{\text{tar}}(c_{i}|\mathcal{C}_{\text{gen}},c_{<i}),$
where $\mathcal{C}_{\text{gen}}$ denotes the input prompt constructed using $\{I_{\text{gen}},x\}$ [^2] We provide the prompt templates for evidence compression and context generation in Table 10. and $|c|$ denotes the length of the generated context $c$.
@@
$c_{i}$ $=\arg\max_{c^{\prime}_{i},c^{\prime\prime}_{i}\in V}(\alpha\cdot\log P_{\text{tar}}(c^{\prime}_{i}\mid\mathcal{C}_{\text{gen}},c_{<i})$
$+(1-\alpha)\cdot\log P_{\text{comp}}(c^{\prime\prime}_{i}\mid\mathcal{C}_{\text{comp}},c_{<i})),$
where $c_{i}$ is the subsequent token, and $\alpha$ is the ensemble coefficient that weighs between the two probability distributions. We demonstrate how the coefficient $\alpha$ impacts both the perplexity and the downstream performance in Section˜4.2.
Ensemble decoding proactively shifts the token search space in evidence compression by upweighting those tokens with lower perplexity from the target model’s perspective, resulting in a compressed evidence that is more familiar to the target model. Note that since both objectives ultimately share the goal of generating context relevant to the question, combining the logits ensures alignment with this ultimate goal.
In addition, ensemble decoding enables FaviComp to seamlessly integrate both retrieval knowledge from the external evidence set and the target model’s parametric knowledge. Specifically, FaviComp selects the $\arg\max$ token from the target model only when the token’s probability is higher than that of the compression model, demonstrating that FaviComp draws on parametric knowledge only when necessary—potentially when the compression model is uncertain about the next token. This is particularly beneficial for complex tasks like multi-document QA, where the evidence set may not include all the necessary information (Mallen et al., 2023). In such cases, the missing information in compressed evidence can be supplemented by tokens generated from context generation by the target model, which is entirely based on parametric knowledge. We demonstrate in Section˜4.3 and Section˜5 that FaviComp can incorporate knowledge from both sources effectively, leading to a performance boost compared to compression methods that solely focus on distilling knowledge from the evidence set.
@@ 3.2 Implementation Details
For all the comparison methods, we utilize Llama3-8B-Instruct and Mixtral-8x7B- Instruct as the target model to tackle downstream QA tasks with RAG. For FaviComp and Zero-shot Summarization, we employ two compression models, one for each target model: Llama3.2-3B-Instruct for Llama3- 8B-Instruct target model and Mistral-7B- Instruct for Mixtral-8x7B-Instruct target model. For each question, we retrieve five documents from 2018 Wikipedia corpus (Karpukhin et al., 2020) using Contriever-MSMARCO (Izacard et al., 2021), so as to be consistent with previous studies (Xu et al., 2024; Yoon et al., 2024). We set ensemble coefficient $\alpha$ of FaviComp to 0.5 by default, for which more analyses are given in Section˜4.2. The prompts used in the experiment are presented in Appendix˜C.
@@ 3.3 Baselines
We consider the following categories of baselines. (1) No Context: RAG without any context. (2) Gold Compression: RAG using directly relevant evidence from the retrieved documents if they exist. (3) Raw Document: RAG with raw documents that have not undergone any compression. (4) Generated Context (Yu et al., 2023): RAG with context generated by the same LM as the target model. This is equivalent to FaviComp with $\alpha=1$, as we rely solely on the target model to generate context when $\alpha=1$. (5) Reranking-based Methods: We rerank sentences in the evidence set and choose top-ranked sentences as the context. We utilize two rerankers—Sentence-BERT (Reimers and Gurevych, 2020) and RECOMP-extractive (Xu et al., 2024). (6) Compression-based Methods: We employ four compressors—LongLLMLingua (Jiang et al., 2023a), RECOMP-abstractive (Xu et al., 2024), CompAct (Yoon et al., 2024), and Zero-shot Summarization. For Zero-shot Summarization, we use the same evidence compression instruction prompt of FaviComp to summarize multiple evidence using the same LM as the target model. This is equivalent to FaviComp with $\alpha=0$, as we depend entirely on the compression model without any intervention from the target model. [^3] A more detailed explanation of the implementation of the baselines is provided in Appendix A.
@@
Table 1: Experimental results on five open-domain QA datasets. Size column represents the size of the compression model used for each method. indicates a fully-supervised compression model, where the compressor is trained.
@@ 4 Experimental Results
In this section, we compare the overall performance of FaviComp with other baselines across the five datasets (Section˜4.1), explore the impact of ensemble coefficient $\alpha$ on performance and perplexity (Section˜4.2), investigate how effectively FaviComp incorporate parametric and non-parametric knowledge (Section˜4.3), and compare the compression rates with other baselines (Section˜4.4).
@@ 4.1 Main Results
The overall performance of FaviComp and the baselines across the five datasets are presented in Table˜1.444We present additional experimental results using other combinations of compression and target model at Section B.1. To start with, the compression-based methods consistently outperform the reranking-based methods, due to the fact that the reranking-based methods are prone to losing more question-relevant information by discarding lower-ranked sentences.
@@
Finally, given that Zero-shot Summarization corresponds to FaviComp with $\alpha=0$ and Generated Context corresponds to FaviComp with $\alpha=1$, the fact that FaviComp outperforms both baselines highlights its ability to effectively incorporate tokens from both sources—evidence summary and generated context. This results in superior performance compared to relying on one source alone.
@@ 4.2 Impact of Ensemble Coefficient on Performance and Perplexity
Figure˜2 illustrates how performance and perplexity change as the ensemble coefficient $\alpha$ is varied across the values when using Llama3.2-3B-Instruct and Llama3-8B- Instruct compression-target pairs on HQA and MQ datasets [^6] Results for other datasets are included in Figure 6.. We calculate the perplexity of the compressed evidence conditioned on the preceding inputs, i.e. instruction, demonstrations, and the question. For all the datasets, performance is the highest when $\alpha=0.5$, indicating that proactively lowering perplexity by equally weighting both input sources yields the best results. When $\alpha$ is below 0.5, performance improves as the perplexity of compressed evidence decreases, which aligns with the previous works (Liu et al., 2024; Gonen et al., 2023). However, when $\alpha$ exceeds 0.5, performance declines as perplexity decreases due to the lack of evidential knowledge during evidence compression. Additionally, when $\alpha$ reaches 0.9 or 1.0, there is a slight rise in the perplexity due to LM’s increased uncertainty with limited evidential knowledge.
@@
Methods NQ TQA HQA Wiki MQ Generated Context 36.6 62.4 29.7 28.7 14.8 Zero-shot Summarization 43.2 67.1 38.5 31.1 15.3 Concatenation 42.5 66.7 36.5 29.0 15.6 FaviComp 46.8 70.9 41.6 35.2 19.9
Table 2: Performance (F1) comparison against concatenation of parametric and non-parametric knowledge.
4.3 Integration of Parametric and Non-parametric Knowledge
As shown in Figure˜3, we compare the accuracy of FaviComp with Llama3.2-3B-Instruct and Llama3-8B-Instruct compression-target pairs on $\mathrm{Hits=0}$ and $\mathrm{Hits=1}$ subsets with the top-performing baselines, Zero-shot Summarization and CompAct [^7] We provide results of FaviComp on various alpha values in Section B.3. FaviComp outperforms other baselines in the $\mathrm{Hits=0}$ subset while performing comparably with others in the $\mathrm{Hits=1}$ subset. This proves that FaviComp effectively relies on parametric knowledge rather than evidential knowledge when faced with irrelevant evidence, while maintaining similar effectiveness in utilizing evidential knowledge when relevant evidence is present.
In addition, we conduct another experiment to demonstrate FaviComp’s superior ability to synergize two sources of knowledge. We compare it against a straightforward approach that concatenates parametric and non-parametric knowledge as context for downstream generation. Specifically, we concatenate the compressed evidence from the Zero-shot Summarization with the generated context from the Generated Context and use this concatenated context for evaluation. The results, shown in Table˜2, reveal that simple concatenation underperforms compared to the Zero-shot Summarization baseline. This suggests that naively merging non-parametric and parametric knowledge in-context can be less effective than relying solely on non-parametric knowledge. In contrast, FaviComp effectively integrates both knowledge sources during compression, leveraging their synergy to achieve superior performance.
4.4 Compression Rate Comparisons
[…]